# Knowledge Transfer from Simple to Complex: A Safe and Efficient Reinforcement Learning Framework for Autonomous Driving Decision-Making

### v3

@@ 3.1 Proximal Policy Optimization Algorithm
Eq. 7 contains $\nu^{\pi_{\theta^{\prime}}}(s)$, which results from the interaction between the new policy $\pi_{\theta^{\prime}}$ and the environment. However, $\pi_{\theta^{\prime}}$ is the exact policy we need to solve. To address this, we replace $\nu^{\pi_{\theta}}(s)$ with $\nu^{\pi_{\theta^{\prime}}}(s)$ in the total discounted rewards function $J(\theta)$. Thus, the local approximate objective function is obtained:
$\centering L_{\theta}(\theta^{\prime})&=J(\theta)+\dfrac{1}{1-\gamma}\sum_{s}\nu^{\pi_{\theta}}(s)\sum_{a}\pi_{\theta^{\prime}}(a|s)A^{\pi_{\theta}}(s,a)\\ &=J(\theta)+\dfrac{1}{1-\gamma}\mathbb{E}_{s\sim\nu^{\pi_{\theta}}}\mathbb{E}_{a\sim\pi_{\theta^{\prime}}(\cdot|s_{t})}[A^{\pi_{\theta}}(s,a)]\end{split}$ (8)
Since $J(\theta)=\mathbb{E}_{\pi_{\theta}}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\right]$, only the second term in Eq. 8 is related to optimizing the policy $\pi_{\theta^{\prime}}$. We employ the importance sampling method to handle the action distribution, resulting in the following surrogate optimization objective function:
$\centering L_{\theta}(\theta^{\prime})&=\mathbb{E}_{s\sim\nu^{\pi_{\theta}}}\mathbb{E}_{a\sim\pi_{\theta^{\prime}}(\cdot|s_{t})}[A^{\pi_{\theta}}(s,a)]\\ &=\mathbb{E}_{s\sim\nu^{\pi_{\theta}}}\mathbb{E}_{a\sim\pi_{\theta}(\cdot|s_{t})}\left[\dfrac{\pi_{\theta^{\prime}}(a|s)}{\pi_{\theta}(a|s)}A^{\pi_{\theta}}(s,a)\right]\end{split}$ (9)
Hence, we can estimate and optimize the new policy $\pi_{\theta^{\prime}}$ using data sampled from the old policy $\pi_{\theta}$. To ensure that the objective function meets the necessary conditions, Schulman et al. (2015) derived the following performance bounds:
Theorem 1.
Define that $C=4\gamma\max_{s,a}|A^{\pi_{\theta}}(s,a)|/(1-\gamma)^{2}$, $D^{s}_{KL}(\pi_{\theta},\pi_{\theta^{\prime}})\stackrel{{\scriptstyle\triangle}}{{=}}D_{KL}(\pi_{\theta}(\cdot|s)\parallel\pi_{\theta^{\prime}}(\cdot|s))$, $D^{max}_{KL}(\pi_{\theta},\pi_{\theta^{\prime}})\stackrel{{\scriptstyle\triangle}}{{=}}\max_{s\in S}D^{s}_{KL}(\pi_{\theta},\pi_{\theta^{\prime}})$, $M_{\theta}(\theta^{\prime})=L_{\theta}(\theta^{\prime})-CD^{max}_{KL}(\pi_{\theta},\pi_{\theta^{\prime}})$. We have $J(\theta^{\prime})\geq M_{\theta}(\theta^{\prime})$, $J(\theta)=M_{\theta}(\theta)$.
This theorem indicates that by maximizing $M_{\theta}(\theta^{\prime})$, the performance of the new policy $\pi_{\theta^{\prime}}$ is guaranteed to be non-decreasing. However, due to the deep architecture of the policy network, using an excessively large step size when optimizing $L_{\theta}(\theta^{\prime})$ along the policy gradient can lead to substantial changes in the policy, potentially resulting in overshooting and negatively impacting performance. To address this issue, Schulman et al. (2015) proposed Trust Region Policy Optimization (TRPO), which optimizes $L_{\theta}(\theta^{\prime})$ with a constraint on the KL divergence, $D^{max}_{KL}(\pi_{\theta},\pi_{\theta^{\prime}})\leq\delta$, between the old and new policies. To simplify the computation, the average KL divergence can be used to relax the constraint on the maximum value of the KL divergence. The final simplified surrogate optimization objective for the TRPO algorithm is:
@@
However, TRPO is computationally complex and inefficient. To address this, Schulman et al. (2017) proposed the PPO algorithm, which replaces the KL divergence constraint in TRPO with a direct maximization of the clipped surrogate objective function:
@@
where the probability ratio $r_{t}(\theta)=\dfrac{\pi_{\theta^{\prime}}(a_{t}|s_{t})}{\pi_{\theta}(a_{t}|s_{t})}$ is clipped within the range $(1-\epsilon,1+\epsilon)$ using the operator $clip$, defined as $clip(r_{t}(\theta),1-\epsilon,1+\epsilon)=\min(\max(r_{t}(\theta),1-\epsilon),1+\epsilon)$, where $\epsilon$ is a hyperparameter that represents the clipping range. The minimization operation ensures that Eq. 11 serves as the lower bound of Eq. 9.
3.2 Adaptive Clipping Proximal Policy Optimization Algorithm
Although the clipping mechanism in the PPO algorithm provides an effective and reliable strategy for policy updates, a fixed clipping factor $\epsilon$ may not adaptively improve learning performance based on the significance of each sample, particularly when the samples are come from different sources. To address this issue, we propose an improved PPO algorithm called ACPPO. This algorithm incorporates an adaptive clipping mechanism, allowing for automatic adjustment of the clipping factor based on the importance of the samples. When encountering more significant samples, the algorithm permits larger updates to the policy parameters, while constraining changes to the policy parameters for less critical samples.
First, we analyze the clipping mechanism of the PPO algorithm and establish its relationship with the objective constraints. Based on Eq. 7 and Eq. 9, we derive the following:
$\begin{split}J(\theta^{\prime})-J(\theta)&=\dfrac{1}{1-\gamma}\mathbb{E}_{s\sim\nu^{\pi_{\theta^{\prime}}}}\mathbb{E}_{a\sim\pi_{\theta^{\prime}}(\cdot|s_{t})}[A^{\pi_{\theta}}(s,a)]\\ &=\dfrac{1}{1-\gamma}L_{\theta}(\theta^{\prime})+\dfrac{1}{1-\gamma}\left[\mathbb{E}_{s\sim\nu^{\pi_{\theta^{\prime}}}}\mathbb{E}_{a\sim\pi_{\theta^{\prime}}(\cdot|s_{t})}[A^{\pi_{\theta}}(s,a)]-\mathbb{E}_{s\sim\nu^{\pi_{\theta}}}\mathbb{E}_{a\sim\pi_{\theta^{\prime}}(\cdot|s_{t})}[A^{\pi_{\theta}}(s,a)]\right]\\ &\geq\dfrac{1}{1-\gamma}L_{\theta}(\theta^{\prime})-\dfrac{1}{1-\gamma}\left|\sum_{s,a}\left[\nu^{\pi_{\theta^{\prime}}}(s)-\nu^{\pi_{\theta}}(s)\right]\times\pi_{\theta^{\prime}}(a|s)A^{\pi_{\theta}}(s,a)\right|\\ &=\dfrac{1}{1-\gamma}L_{\theta}(\theta^{\prime})-\dfrac{1}{1-\gamma}\left|\sum_{s}\left[\nu^{\pi_{\theta^{\prime}}}(s)-\nu^{\pi_{\theta}}(s)\right]\times\sum_{a}\pi_{\theta^{\prime}}(a|s)A^{\pi_{\theta}}(s,a)\right|\end{split}$ (12)
We define that $\Delta\nu^{\pi_{\theta^{\prime}}}_{\pi_{\theta}}\cdot A^{\pi_{\theta^{\prime}}}_{\pi_{\theta}}\stackrel{{\scriptstyle\triangle}}{{=}}\sum_{s}\left[\nu^{\pi_{\theta^{\prime}}}(s)-\nu^{\pi_{\theta}}(s)\right]\times\sum_{a}\pi_{\theta^{\prime}}(a|s)A^{\pi_{\theta}}(s,a)$, and bound it by Hölder’s inequality:
$\begin{split}\lvert\Delta\nu^{\pi_{\theta^{\prime}}}_{\pi_{\theta}}\cdot A^{\pi_{\theta^{\prime}}}_{\pi_{\theta}}\rvert&\leq\sum_{s}\left[|\nu^{\pi_{\theta^{\prime}}}(s)-\nu^{\pi_{\theta}}(s)|\right]\max_{s,a}\left|A^{\pi_{\theta}}(s,a)\right|\\ &\leq\dfrac{\gamma}{1-\gamma}\max_{s,a}\left|A^{\pi_{\theta}}(s,a)\right|\sum_{s}\nu^{\pi_{\theta}}(s)\sum_{a}\left|\pi_{\theta^{\prime}}(a|s)-\pi_{\theta}(a|s)\right|\\ &=\dfrac{\gamma}{1-\gamma}\max_{s,a}\left|A^{\pi_{\theta}}(s,a)\right|\mathbb{E}_{s\sim\nu^{\pi_{\theta}}}\mathbb{E}_{a\sim\pi_{\theta}(\cdot|s_{t})}\left[\dfrac{\pi_{\theta^{\prime}}(a|s)}{\pi_{\theta}(a|s)}-1\right]\\ &=\dfrac{\gamma}{1-\gamma}\max_{s,a}\left|A^{\pi_{\theta}}(s,a)\right|\mathbb{E}_{s\sim\nu^{\pi_{\theta}}}\mathbb{E}_{a\sim\pi_{\theta}(\cdot|s_{t})}\left[r_{t}(\theta)-1\right]\end{split}$ (13)
Thus, we can get the following improvement lower bound for policy as stated in Theorem 2:
Theorem 2.
For the current policy $\pi_{\theta^{\prime}}$ and the old policy $\pi_{\theta}$, we have:
$J(\theta^{\prime})-J(\theta)\geq\dfrac{1}{1-\gamma}L_{\theta}(\theta^{\prime})-C\cdot\mathbb{E}_{s\sim\nu^{\pi_{\theta}}}\mathbb{E}_{a\sim\pi_{\theta}(\cdot|s_{t})}\left|r_{t}(\theta)-1\right|$ (14)
where $C$ is defined as in Theorem 1.
Theorem 2 demonstrates that the performance deviation between any two successive policies can be constrained by the expected value $\mathbb{E}_{s\sim\nu^{\pi_{\theta}}}\mathbb{E}_{a\sim\pi_{\theta}(\cdot|s_{t})}\left|r_{t}(\theta)-1\right|$, which represents the absolute value of the policy probability deviation. According to Eq. 11, the gradient of the clipped surrogate objective is given by:
$\nabla_{\theta}L^{CLIP}(\theta)=\mathbb{E}_{\pi_{\theta}}\left[r_{t}(\theta)A^{\pi_{\theta}}(s,a)\nabla_{\theta}\log\pi_{\theta}(a_{t}|s_{t})\times\left\{r_{t}(\theta)\in(1-\epsilon,1+\epsilon)\text{or}\text{sgn}(r_{t}(\theta)-1)\neq\text{sgn}(A^{\pi_{\theta}}(s,a))\right\}\right]$ (15)
PPO limits the difference between successive policies by eliminating the reward generated when $\left|r_{t}(\theta)-1\right|$ exceeds the clipping parameter $\epsilon$, while maximizing the surrogate objective. This is achieved by using the clip operator to restrict the values of $r_{t}(\theta)$ within a fixed range $(1-\epsilon,1+\epsilon)$, thereby limiting the variation between the new and old policies and ensuring that the variance among samples does not become excessive. The value of the objective function $L^{CLIP}(\theta)$ depends on whether $A^{\pi_{\theta}}$ is positive or negative. If $A^{\pi_{\theta}}>0$, indicating that the state-action value of $a_{t}$ is above the average state value, maximizing $L^{CLIP}(\theta)$ increases $r_{t}(\theta)$, but restricts it to remain below $1+\epsilon$. Conversely, if $A^{\pi_{\theta}}<0$, meaning the state-action value of $a_{t}$ is below the average state value, maximizing $L^{CLIP}(\theta)$ decreases $r_{t}(\theta)$ , but restricts it to remain above $1-\epsilon$.
From the aforementioned conclusions, it is clear that the use of collected samples to update policies is constrained by the clipping factor $\epsilon$. In traditional PPO algorithms, $\epsilon$ is fixed. When encountering samples from different sources, a fixed clipping factor may fail to fully leverage the information from these diverse samples, especially for important ones, limiting the scope of policy updates and resulting in reduced algorithm performance or slower convergence rates. When the student agent interacts with the environment and receives direct interventions on actions from teacher model, the actions collected by the agent in the same state $s_{t}$ originate from two different sources:
  1. 1.
    The action $a^{s}_{t}$ output by the student policy;
  2. 2.
    The action $a^{t}_{t}$ output by the teacher policy.
Assuming that the teacher’s actions $a^{t}_{t}$ are safe and reliable in most cases, we can consider that $a^{t}_{t}$ has a greater advantage over $a^{s}_{t}$ for the current state $s_{t}$. Therefore, $a^{t}_{t}$ is more important for agent training. To accommodate samples from different sources, we propose using a dynamically adjusted adaptive clipping factor to better reflect the importance of the samples. The following adjustments are made to the surrogate objective function:
@@ 3.2 Adaptive Clipping Proximal Policy Optimization Algorithm
$L^{ACLIP}(\theta)=\left\{\begin{array}[]{lr}\mathbb{E}_{\pi_{\theta}}\left[\min\left(r_{t}(\theta)A^{\pi_{\theta}}(s,a),\text{clip}\left(r_{t}(\theta),1-(\epsilon+\epsilon^{\prime}),1+(\epsilon-\epsilon^{\prime})\right)A^{\pi_{\theta}}(s,a)\right)\right],&{if}\ a_{t}=a^{s}_{t}\\ \mathbb{E}_{\pi_{\theta}}\left[\min\left(r_{t}(\theta)A^{\pi_{\theta}}(s,a),\text{clip}\left(r_{t}(\theta),1-(\epsilon-\epsilon^{\prime}),1+(\epsilon+\epsilon^{\prime})\right)A^{\pi_{\theta}}(s,a)\right)\right],&{if}\ a_{t}=a^{t}_{t}\\ \end{array}\right.$ (16)
where $\epsilon^{\prime}$ is the adaptive factor, determined by the following equation:
@@
where $\psi$ is a hyperparameter used to adjust the extent of adaptive clipping, and $\pi_{\theta}(a^{s}_{t}|s_{t})-\pi_{\theta}(a^{t}_{t}|s_{t})$ represents the difference in the probabilities of policy $\pi_{\theta}$ outputting $a^{s}_{t}$ and $a^{t}_{t}$ in state $s_{t}$. It should be noted that, in order to maintain the monotonicity of the objective function, the sign of the adjusted clipping factor must not change, so $(\epsilon-\epsilon^{\prime})\geq 0$.
For a clearer understanding of Eq. 16, we analyze it using Fig. 1. It is evident that when $a_{t}=a^{s}_{t}$, we consider $a^{s}_{t}$ to be less advantageous than $a^{t}_{t}$. Therefore, when $A^{\pi_{\theta}}>0$, we reduce the upper bound of $r_{t}$, decreasing the increment of $\pi_{\theta}(a^{s}_{t}|s_{t})$, and the larger the difference $\pi_{\theta}(a^{s}_{t}|s_{t})-\pi_{\theta}(a^{t}_{t}|s_{t})$, the lower the upper bound of $r_{t}$. Conversely, when $A^{\pi_{\theta}}<0$, we further lower the lower bound of $r_{t}$, making the decrease in $\pi_{\theta}(a^{s}_{t}|s_{t})$ more significant, and the larger the difference $\pi_{\theta}(a^{s}_{t}|s_{t})-\pi_{\theta}(a^{t}_{t}|s_{t})$, the lower the lower bound of $r_{t}$. On the other hand, when $a_{t}=a^{t}_{t}$, we consider this action to be more important. Therefore, when $A^{\pi_{\theta}}>0$, we further raise the upper bound of $r_{t}$, increasing the increment of $\pi_{\theta}(a^{t}_{t}|s_{t})$, and the larger the difference $\pi_{\theta}(a^{s}_{t}|s_{t})-\pi_{\theta}(a^{t}_{t}|s_{t})$, the higher the upper bound of $r_{t}$. When $A^{\pi_{\theta}}<0$, we raise the lower bound of $r_{t}$, reducing the decrement in $\pi_{\theta}(a^{t}_{t}|s_{t})$, and the larger the difference $\pi_{\theta}(a^{s}_{t}|s_{t})-\pi_{\theta}(a^{t}_{t}|s_{t})$, the higher the lower bound of $r_{t}$. Through adaptive clipping, the PPO algorithm can better adapt to the importance of different samples, quickly aligning the policy towards the expert policy during training, thus enhancing the learning efficiency of the agent.
(a) Adaptive clipping when $a_{t}=a^{s}_{t}$
(b) Adaptive clipping when $a_{t}=a^{t}_{t}$
Figure 1: Adaptive clipping range of ACPPO algorithm

### v4

@@ 3.1 Proximal Policy Optimization Algorithm
Eq. 7 includes $\nu^{\pi_{\theta^{\prime}}}(s)$, where $\pi_{\theta^{\prime}}$ is the target policy we aim to solve for. To facilitate the calculation, we substitute $\nu^{\pi_{\theta}}(s)$ for $\nu^{\pi_{\theta^{\prime}}}(s)$ in $J(\theta)$. Given the close similarity between the new and old policies, this approximation is considered acceptable. Thus, we obtain the locally approximated objective function:
$\centering L_{\theta}(\theta^{\prime})=J(\theta)+\dfrac{1}{1-\gamma}\sum_{s}\nu^{\pi_{\theta}}(s)\sum_{a}\pi_{\theta^{\prime}}(a|s)A^{\pi_{\theta}}(s,a)=J(\theta)+\dfrac{1}{1-\gamma}\mathbb{E}_{s\sim\nu^{\pi_{\theta}}}\mathbb{E}_{a\sim\pi_{\theta^{\prime}}(\cdot|s_{t})}[A^{\pi_{\theta}}(s,a)]\end{split}$ (8)
Since $J(\theta)=\mathbb{E}_{\pi_{\theta}}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\right]$, the following objective function can be derived using importance sampling:
$\centering L_{\theta}(\theta^{\prime})&=\mathbb{E}_{s\sim\nu^{\pi_{\theta}}}\mathbb{E}_{a\sim\pi_{\theta^{\prime}}(\cdot|s_{t})}[A^{\pi_{\theta}}(s,a)]=\mathbb{E}_{s\sim\nu^{\pi_{\theta}}}\mathbb{E}_{a\sim\pi_{\theta}(\cdot|s_{t})}\left[\dfrac{\pi_{\theta^{\prime}}(a|s)}{\pi_{\theta}(a|s)}A^{\pi_{\theta}}(s,a)\right]\end{split}$ (9)
Hence, by performing gradient descent on the objective function, the policy $\pi_{\theta}$ can be optimized. However, the deep structure of neural networks can lead to abrupt policy shifts if an inappropriate step size is used when optimizing the objective function along the policy gradient, potentially causing significant performance degradation. Schulman et al. (2015) proposed the Trust Region Policy Optimization (TRPO) method, which optimizes $L_{\theta}(\theta^{\prime})$ by imposing a constraint on the KL divergence between the new and old policies. To simplify the computation, the average KL divergence can be used to relax the constraint on the maximum value of the KL divergence, so the optimization objective for TRPO is:
@@
Although TRPO effectively limits the size of policy updates, it is computationally complex and time-consuming. To address this, the PPO algorithm (Schulman et al., 2017) constrains policy updates by maximizing a clipped objective function:
@@
where $r_{t}(\theta)=\dfrac{\pi_{\theta^{\prime}}(a_{t}|s_{t})}{\pi_{\theta}(a_{t}|s_{t})}$ is the probability ratio, which is clipped within the range $(1-\epsilon,1+\epsilon)$ by the operator $clip$; $\epsilon$ is the clipping factor that controls the clipping range.
3.2 Adaptive Clipping Proximal Policy Optimization Plus Algorithm
According to Eq. 11, the value of $L^{CLIP}(\theta)$ is primarily influenced by the $A^{\pi_{\theta}}$. When $A^{\pi_{\theta}}$ is positive, it indicates that the action $a_{t}$ is better than the average level, making it a favorable choice. In this case, maximizing $L^{CLIP}(\theta)$ along the gradient direction increases $r_{t}(\theta)$, thereby raising the probability of selecting $a_{t}$. Conversely, when $A^{\pi_{\theta}}$ is negative, it implies that $a_{t}$ is below average and thus not an ideal choice. In this situation, maximizing $L^{CLIP}(\theta)$ along the gradient direction decreases $r_{t}(\theta)$, reducing the probability of selecting $a_{t}$. PPO restricts $r_{t}(\theta)$ within the range $(1-\epsilon,1+\epsilon)$ through clipping.
This clipping mechanism limits the magnitude of policy updates in a simple and effective manner, thereby avoiding performance degradation due to abrupt policy changes. However, previous PPO-based algorithms typically set the clipping factor $\epsilon$ as a fixed value. When samples come from different sources with significant importance differences, this fixed value cannot adapt flexibly to the characteristics of each sample, making it challenging to accurately control the clipping range. To address this issue, we proposed the improved ACPPO algorithm in previous work (Zhou et al., 2024). This algorithm introduces an adaptive clipping mechanism that dynamically adjusts $\epsilon$ based on sample importance, using the probability distribution of different actions. The modified objective function is as follows:
@@ 3.2 Adaptive Clipping Proximal Policy Optimization Plus Algorithm
$L^{ACLIP}(\theta)=\left\{\begin{array}[]{lr}\mathbb{E}_{\pi_{\theta}}\left[\min\left(r_{t}(\theta)A^{\pi_{\theta}}(s,a),\text{clip}\left(r_{t}(\theta),1-(\epsilon+\epsilon^{\prime}),1+(\epsilon-\epsilon^{\prime})\right)A^{\pi_{\theta}}(s,a)\right)\right],&{if}\ a_{t}=a^{P}_{t}\\ \mathbb{E}_{\pi_{\theta}}\left[\min\left(r_{t}(\theta)A^{\pi_{\theta}}(s,a),\text{clip}\left(r_{t}(\theta),1-(\epsilon-\epsilon^{\prime}),1+(\epsilon+\epsilon^{\prime})\right)A^{\pi_{\theta}}(s,a)\right)\right],&{if}\ a_{t}=a^{E}_{t}\\ \end{array}\right.$ (12)
where action $a^{E}_{t}$ originates from the expert policy and is considered safe and reliable in most situations; action $a^{P}_{t}$ is generated by the PPO policy and may lead to suboptimal outcomes, especially in the early training stages when the policy network is not yet fully optimized; $\epsilon^{\prime}$ is the adaptive clipping factor, defined as follows:
@@
where $\pi_{\theta}(\cdot|s_{t})$ is the probability distribution of different actions in state $s_{t}$; $\psi$ serves as an adaptive factor to regulate the clipping range.
Although ACPPO performs adaptive clipping based on sample importance, the samples collected at each step are derived solely from either $a^{P}_{t}$ or $a^{E}_{t}$. When the expert policy fully intervenes, it is highly likely that all collected samples are from expert actions, which prevents the agent from learning to avoid incorrect decisions, leading to an imbalance between positive and negative samples. In this study, we collect the state, actions $a^{t}_{t}$ and $a^{s}_{t}$ generated by both the teacher and student, and the corresponding return value at each step as MDP training samples. This approach not only doubles the amount of training data compared to ACPPO, improving sample efficiency, but also collects data for different actions under the same state, effectively mitigating the positive-negative sample imbalance issue in ACPPO. We refer to the improved algorithm as Adaptive Clipping Proximal Policy Optimization Plus (ACPPO+), with its objective function modified as follows:
$\begin{split}L^{ACLIP+}(\theta)&=\mathbb{E}_{\pi_{\theta}}\Bigg{[}\mathbb{I}_{a_{t}^{s}}\cdot\min\left(r_{t}(\theta)A^{\pi_{\theta}}(s,a_{t}^{s}),\text{clip}\left(r_{t}(\theta),1-(\epsilon+\epsilon^{\prime}),1+(\epsilon-\epsilon^{\prime})\right)A^{\pi_{\theta}}(s,a_{t}^{s})\right)\\ &\quad+\mathbb{I}_{a_{t}^{t}}\cdot\min\left(r_{t}(\theta)A^{\pi_{\theta}}(s,a_{t}^{t}),\text{clip}\left(r_{t}(\theta),1-(\epsilon-\epsilon^{\prime}),1+(\epsilon+\epsilon^{\prime})\right)A^{\pi_{\theta}}(s,a_{t}^{t})\right)\Bigg{]}\end{split}$ (14)
where $\mathbb{I}_{a_{t}^{s}}$ and $\mathbb{I}_{a_{t}^{t}}$ are indicator variables representing whether actions $a_{t}^{s}$ and $a_{t}^{t}$ exist, respectively. If the action exists, the value is 1; otherwise, it is 0.
As seen from Eq. 14, ACPPO+ allows for greater updates to policy parameters when training with more important teacher samples, while limiting the influence of student samples. Through this adaptive clipping mechanism, ACPPO+ not only leverages data generated by both teacher and student policies but also effectively adapts to samples from various sources. This accelerates the alignment of the student policy with the teacher policy, thereby improving the agent’s learning efficiency.